OpenStudy (anonymous):

b-(5/b)=4 solve and check solution (i know final answer is -1,5 but i dont know how to get to it pls show work)

OpenStudy (anonymous):

multiply both sides by b, then solve the quadratic

OpenStudy (anonymous):

i did b(b)-b(5/b)=4b got b^2-5=4b but not sure how to continue

OpenStudy (amistre64):

rearragnge into a quadratic: b^2 -4b -5 = 0

OpenStudy (amistre64):

facotr it out; (b-5)(b+1) = 0 b = -1,5

OpenStudy (anonymous):

thx yet agian man XD

OpenStudy (amistre64):

:) yw